# Why Rectified Power Unit Networks Fail and How to Improve It: An Effective Theory Perspective

### v3

@@ Abstract
The Rectified Power Unit (RePU) activation functions, unlike the Rectified Linear Unit (ReLU), have the advantage of being a differentiable function when constructing neural networks. However, it can be experimentally observed when deep layers are stacked, neural networks constructed with RePU encounter critical issues. These issues include the values exploding or vanishing and failure of training. And these happen regardless of the hyperparameter initialization. From the perspective of effective theory, we aim to identify the causes of this phenomenon and propose a new activation function that retains the advantages of RePU while overcoming its drawbacks.
keywords:
Neural Networks , Activation Functions , Deep Learning Theory , Effective Theory , Rectified Power Units
journal: Neural Networks
[…]
organization=Department of Mathematical Sciences, Seoul National University,addressline=1, Gwanak-ro, Gwanak-gu, city=Seoul, postcode=08826, state=Seoul, country=South Korea
@@ 1.1 Activation Functions
An activation function is a function that acts on the output of each layer of a neural network. The remarkable success of deep neural networks is closely related to the choice of appropriate nonlinear activation functions, and there has been extensive discussion solely on the research of activation functions. Early studies on neural networks include the perceptron [1], where the activation function is the sign function, and it was composed of only one or two layers. The perceptron had the drawback that its derivative was almost everywhere zero, making it difficult to apply the backpropagation algorithm. According to Cybenko [2], for a neural network to satisfy the universal approximation property, the activation function must be bounded, non-constant, monotonically increasing, and continuous. In the 1980s and 1990s, smooth activations like the sigmoid activation and tanh activation, along with their variations, were proposed [3], [4], [19]. These activations had the disadvantage of gradient vanishing when stacking layers deeply. To overcome this drawback, the Rectified Linear Unit (ReLU) was proposed by Nair and Hinton [5], whose derivative is a non-zero constant. Variations of ReLU, such as PReLU, GELU, and ELU, have also been proposed in the 2010s [7], [8], [9]. Recently, there have been attempts to find activations tailored to specific tasks to impart inductive bias to the neural network structure according to the task [10], [11].
@@ 1.2 Rectified Power Unit
One of the attempts to provide good regularity and inductive bias to the architecture is the Rectified Power Unit (RePU), a generalization of ReLU [12]. Early results using RePU include the application of ReQU, a type of RePU, in the DeepRitz method, which is a neural network approach for solving Partial Differential Equations (PDEs) [13]. RePU has also been proposed to construct differentiable neural networks, and error bounds related to this have been analyzed [12]. According to research, using a RePU network allows for the optimal approximation of smooth functions in Sobolev space with optimal depth and width, and the optimal approximation error bound is also provided in [12]. Similarly, another study derived that a shallow RePU network can approximate multivariate functions of a certain function class with finite neural network weight norms and possibly unbounded width [14]. Furthermore, it was shown that the derivatives of a network composed of RePU can be expressed as a mixture of various types of RePU, and an improved upper bound on the complexity of these derivatives was derived. This demonstrates the capability of RePU by showing the risk bound for the deep score matching estimator (DSME) [15]. Despite these results highlighting the good regularity of RePU neural networks, experimental evidence shows that RePU networks suffer from exploding or vanishing values and fail to train properly when the layers are deep.
1.3 Effective Theory of Neural Networks
The language of quantum field theory (QFT) is useful for analyzing systems with a very high degree of freedom. One such example is artificial neural networks. It is known that the distribution of neural network ensembles follows a Gaussian process when the width is infinite, corresponding to the free field in the QFT framework [16], [17], [18], [24], [25]. In practice, neural networks with finite width exhibit non-Gaussianity in their distributions as layers are stacked, allowing the distribution of neurons to be calculated perturbatively, as in weakly-interacting theory [24], [25]. Interestingly, when the width of the neural network is finite, the propagation of information through the neural network can be understood as a renormalization group (RG) flow. By analyzing the recursive relations of the kernels in RG flow, hyperparameters can be tuned to a criticality state that ensures the stability of the network [18]. Additionally, this analysis allows for the classification of activation functions into universality classes based on the behavior of the RG flow [18]. Some types of activation functions, according to this classification, cannot satisfy criticality conditions from the outset.
@@ 1.4 Our Contribution
We theoretically analyze the RePU activation from the perspective of effective theory. Specifically, we calculate the susceptibility of the RePU activation to theoretically demonstrate that RePU cannot satisfy the criticality condition. We then verify this experimentally by calculating the empirical kernel of the initialized neural network and during the training process. Not only do we identify the fundamental failure of RePU, but we also propose a modified RePU (MRePU) based on theoretical predictions to address these issues. We demonstrate the improvements experimentally by evaluating the empirical kernel and testing performance of MRePU, as compared to RePU.
@@ 2 Preliminary
In this section, we define the neural network and activation functions that will be the subject of our analysis, focusing particularly on the Rectified Power Unit (RePU). Additionally, we explore the effective theory of neural networks, which will be the primary tool for our analysis.
@@ 2.1 Overview of Neural Networks and Activation Functions
Definition 1 (Fully Connected Network (FCN)) A Fully Connected Network $f(x;\theta):\mathbb{R}^{n_{in}}\rightarrow\mathbb{R}^{n_{out}}$ is defined by the following recursive equations:
@@
Definition 2 (Initialization distribution of biases and weights) In the definition of neural networks, there are adjustable weights and biases parameters. The method of setting these parameters at the beginning of training is called initialization, and typically, each parameter follows a specific probability distribution, referred to as the initialization distribution. Although it is possible to use mathematically complex initialization distributions, it is common practice to assign each weight and bias under the condition of being independent and identically distributed (i.i.d). The probability distributions used can include Gaussian distribution, truncated Gaussian distribution, uniform distribution, and others. For wide neural networks, it is known that when the mean of the weight distribution is 0 and the variance is constant, differences arising from different types of distributions are suppressed by $\frac{1}{width}$. Therefore, for the sake of convenience in our discussion, we will set the distributions of weights and biases to follow a mean-zero Gaussian distribution:
@@
Perceptron $\sigma(z)=\begin{cases}1\text{,}&\quad\text{if }z\geq 0\\ 0\text{,}&\quad\text{if }z<0\\ \end{cases}$ [1]
Sigmoid $\sigma(z)=\frac{1}{1+e^{-z}}$ [2]
Tanh $\sigma(z)=\tanh(z)=\frac{e^{z}-e^{-z}}{e^{z}+e^{-z}}$ [2], [19]
Sin $\sigma(z)=sin(z)$ [20], [21]
ReLU $\sigma(z)=\begin{cases}z\text{,}&\quad\text{if }z\geq 0\\ 0\text{,}&\quad\text{if }z<0\\ \end{cases}$ [5], [6]
Leaky ReLU $\sigma(z)=\begin{cases}z\text{,}&\quad\text{if }z\geq 0\\ \alpha z\text{,}&\quad\text{if }z<0\\ \end{cases}$ [7], [22]
Softplus $\sigma(z)=\log(1+e^{z})$ [23]
SWISH $\sigma(z)=\frac{z}{1+e^{-z}}$ [10]
GELU $\sigma(z)=\Big{[}\frac{1}{2}+\frac{1}{2}\text{erf}\Big{(}\frac{z}{\sqrt{2}}\Big{)}\Big{]}z$ [8]
@@
Definition 3 (Rectified Power Unit (RePU) The RePU activation function is defined by the following equation:
@@
where $p$ is a positive integer that determines the power to which the input $z$ is raised. When $p=1$, it corresponds to ReLU, and thus RePU can be considered a generalized version of ReLU. The cases where $p=2$ and $p=3$ are referred to as Rectified Quadratic Unit (ReQU) and Rectified Cubic Unit (ReCU), respectively. For RePU with $p=k$, it is easy to see that it is differentiable $k-1$ times, and this property ensures that the neural network and its gradients are differentiable functions.
@@ 2.2 Effective Field Theory for Neural Networks
Definition 4 ($M$-point correlators) The moments or $M$-point correlators of a probability distribution $p(z)=p(z_{1},...,z_{N})$ are defined as follows:
@@
Definition 5 ($M$-point connected correlators) For $2\leq M$, the cumulant or $M$-point connected correlators of a probability distribution $p(z)$ are defined as follows:
For $M=1$, the connected correlator coincides with the regular correlator. For parity-symmetric distributions, it also coincides for $M=2$. According to Wick’s theorem, it is known that for a Gaussian distribution, the connected correlators for $M>2$ are zero. Therefore, higher-point connected correlators can be used to determine how much the distribution deviates from a Gaussian distribution. Additionally, if the connected correlators for $M>2$ are small, the distribution is defined as nearly-Gaussian. In quantum field theory, the 2-point connected correlator can be seen as an expression of the translational symmetry of a free field. In a free field, there is no interaction due to potential, so all connected correlators higher than the 2-point vanish. Conversely, if there are interactions between particles due to a potential, higher-order connected correlators become non-zero. Similarly, in the neural network we analyze, if the higher-order correlators do not vanish and instead show significant values, this implies meaningful interactions between nodes in the feature space, suggesting that feature learning is taking place.
Induced distributions Following the notation in (1), let $p(z^{(l)}|\mathcal{D})$ denote the probability distribution of the preactivations in the $l$-th layer given a dataset $\mathcal{D}=\{x_{i;\alpha}\}_{i=1,...,n_{0};\alpha=1,...,N_{\mathcal{D}}}$. For the first layer, the preactivation has $M>2$ connected correlators equal to 0, and has the following mean and covariance, indicating that each preactivation follows an independent Gaussian distribution.
$\begin{split}\mathbb{E}[z_{i;\alpha}^{(1)}]&=\mathbb{E}\bigg{[}b_{i}^{(1)}+\sum_{j=0}^{n_{0}}W_{ij}^{(1)}x_{j;\alpha}\bigg{]}=0,\\ \mathbb{E}[z_{i_{1};\alpha_{1}}^{(1)}z_{i_{2};\alpha_{2}}^{(1)}]&=\mathbb{E}\bigg{[}\bigg{(}b_{i_{1}}^{(1)}+\sum_{j=0}^{n_{0}}W_{i_{1}j_{1}}^{(1)}x_{j_{1};\alpha_{1}}\bigg{)}\bigg{(}b_{i_{2}}^{(1)}+\sum_{j=0}^{n_{0}}W_{i_{2}j_{2}}^{(1)}x_{j_{2};\alpha_{2}}\bigg{)}\bigg{]}\\ &=\delta_{i_{1}i_{2]}}\bigg{(}C_{b}^{(1)}+C_{W}^{(1)}\frac{1}{n_{0}}\sum_{j=1}^{n_{0}}x_{j;\alpha_{1}}x_{j;\alpha_{2}}\bigg{)}=\delta_{i_{1}i_{2}}G_{\alpha_{1}\alpha_{2}}^{(1)}.\end{split}$
Where $G_{\alpha_{1}\alpha_{2}}^{(1)}$ is called as the metric, which represents the two-point correlator of the preactivations for different samples. Similarly, the conditional probability $p(z^{(l+1)}|z^{(l)})$ is also described by Gaussian distribution with covariance $\hat{G}_{\alpha_{1}\alpha_{2}}^{(l+1)}$, where the distribution (i.e.$\hat{G}_{\alpha_{1}\alpha_{2}}^{(l+1)}$) depends on the condition $z^{(l)}$ We refer to this $\hat{G}_{\alpha_{1}\alpha_{2}}^{(l+1)}$ as the $l+1$-th layer stochastic metric. We can consider the mean metric $G_{\alpha_{1}\alpha_{2}}^{(l+1)}$, which is the average of the stochastic metric. Additionally, we can consider the variance of $\Delta\hat{G}_{\alpha_{1}\alpha_{2}}^{(l+1)}=\hat{G}_{\alpha_{1}\alpha_{2}}^{(l+1)}-G_{\alpha_{1}\alpha_{2}}^{(l+1)}$, which is the fluctuation of the stochastic metric around the mean metric. We refer to scaled quantity of this variance as the four-point vertex, defined as follows:
@@
$S:=\frac{1}{2}\sum_{i=1}^{n_{l}}\sum_{\alpha_{1}\alpha_{2}\in\mathcal{D}}g^{\alpha_{1}\alpha_{2}}_{(l)}z_{i;\alpha_{1}}^{(l)}z_{i;\alpha_{2}}^{(l)}-\frac{1}{8}\sum_{i_{1},i_{2}=1}^{n_{l}}\sum_{\alpha_{1},...,\alpha_{4}\in\mathcal{D}}v^{(\alpha_{1}\alpha_{2})(\alpha_{3}\alpha_{4})}_{(l)}z_{i;\alpha_{1}}^{(l)}z_{i;\alpha_{2}}^{(l)}z_{i;\alpha_{3}}^{(l)}z_{i;\alpha_{4}}^{(l)}+\cdots.$
It is known that the following relationship holds for the coefficients. This can be derived by comparing the coefficients of the connected correlator given by the mean metric and four-point vertex with those of the connected correlator assumed by the ansatz provided by action $S$. $g^{\alpha_{1}\alpha_{2}}_{(l+1)}$ and $v^{(\alpha_{1}\alpha_{2})(\alpha_{3}\alpha_{4})}_{(l+1)}$ [18]:
$\begin{split}g^{\alpha_{1}\alpha_{2}}_{(l)}&=G^{\alpha_{1}\alpha_{2}}_{(l)}+O(\frac{1}{n_{l-1}}),\\ v^{(\alpha_{1}\alpha_{2})(\alpha_{3}\alpha_{4})}_{(l)}&=\frac{1}{n_{l}}V^{(\alpha_{1}\alpha_{2})(\alpha_{3}\alpha_{4})}_{(l)}+O(\frac{1}{n_{l-1}^{2}}).\end{split}$
@@
Kernels and Representation Group flow Now, we will investigate how the statistics of the preactivation change as they pass through the layers. We will refer to the evolution of the distribution through the layers as the representation group flow, following the terminology of [18]. While it is not feasible to explicitly calculate the statistics for a complex non-Gaussian distribution, it is possible to develop a perturbative theory to understand the recursive flow of the key statistics of the distribution. Before delving into this, let’s define the notation for Gaussian expectations. For an observable $\mathcal{O}(z)$ with a mean-zero Gaussian distribution with variance $\bf{K}=(K_{ij})$, the expectation is denoted as $\langle\mathcal{O}(z)\rangle_{\bf{K}}$. Now, if we define the kernel of the neural network through the following recursive relation:
$\begin{split}K^{(1)}_{\alpha_{1}\alpha_{2}}&:=G^{(1)}_{\alpha_{1}\alpha_{2}},\\ K^{(l+1)}_{\alpha_{1}\alpha_{2}}&=C^{(l+1)}_{b}+C^{(l+1)}_{W}\langle\sigma(z_{\alpha_{1}})\sigma(z_{\alpha_{2}})\rangle_{K^{(l)}}.\end{split}$
it is known that this kernel satisfies following properties, as discussed in [18].
$\begin{split}G^{(l)}_{\alpha_{1}\alpha_{2}}&=K^{(l)}_{\alpha_{1}\alpha_{2}}+O(\frac{1}{n_{l-1}}),\\ V^{(l+1)}_{(\alpha_{1}\alpha_{2})(\alpha_{3}\alpha_{4})}&=\Big{(}C_{W}^{(l+1)}\Big{)}^{2}\Big{[}\langle\sigma(z_{\alpha_{1}})\sigma(z_{\alpha_{2}})\sigma(z_{\alpha_{3}})\sigma(z_{\alpha_{4}})\rangle_{K^{(l)}}\\ &-\langle\sigma(z_{\alpha_{1}})\sigma(z_{\alpha_{2}})\rangle_{K^{(l)}}-\langle\sigma(z_{\alpha_{3}})\sigma(z_{\alpha_{4}})\rangle_{K^{(l)}}\\ &+\frac{n_{l}}{4n_{l-1}}\sum_{\beta_{1},...,\beta_{4}\in\mathcal{D}}V^{(\alpha_{1}\alpha_{2})(\alpha_{3}\alpha_{4})}_{(l)}\Big{\langle}\sigma(z_{\alpha_{1}})\sigma(z_{\alpha_{2}})\Big{(}z_{\beta_{1}}z_{\beta_{2}}-K_{\beta_{1}\beta_{2}}^{(l)}\Big{)}\Big{\rangle}_{K^{(l)}}\\ &\times\Big{\langle}\sigma(z_{\alpha_{3}})\sigma(z_{\alpha_{4}})\Big{(}z_{\beta_{3}}z_{\beta_{4}}-K_{\beta_{3}\beta_{4}}^{(l)}\Big{)}\Big{\rangle}_{K^{(l)}}\Big{]}+O(\frac{1}{n_{l}}).\end{split}$ (4)
@@
Then, it is known that the following recursion can be derived [18]:

### v4

# Why Rectified Power Unit Networks Fail and How to Improve It: An Effective Field Theory Perspective

@@ Abstract
The Rectified Power Unit (RePU) activation function, a differentiable generalization of the Rectified Linear Unit (ReLU), has shown promise in constructing neural networks due to its smoothness properties. However, deep RePU networks often suffer from critical issues such as vanishing or exploding values during training, rendering them unstable regardless of hyperparameter initialization. Leveraging the perspective of effective field theory, we identify the root causes of these failures and propose the Modified Rectified Power Unit (MRePU) activation function. MRePU addresses RePU’s limitations while preserving its advantages, such as differentiability and universal approximation properties. Theoretical analysis demonstrates that MRePU satisfies criticality conditions necessary for stable training, placing it in a distinct universality class. Extensive experiments validate the effectiveness of MRePU, showing significant improvements in training stability and performance across various tasks, including polynomial regression, physics-informed neural networks (PINNs) and real-world vision tasks. Our findings highlight the potential of MRePU as a robust alternative for building deep neural networks.
Keywords: Neural Networks, Activation Functions, Deep Learning Theory, Effective Field Theory, Rectified Power Units
@@ 1.1 Activation Functions
An activation function is a function that acts on the output of each layer of a neural network. The remarkable success of deep neural networks is closely related to the choice of appropriate nonlinear activation functions, and there has been extensive discussion solely on the research of activation functions. Early studies on neural networks include the perceptron (rosenblatt58), where the activation function is the sign function, and it was composed of only one or two layers. The perceptron had the drawback that its derivative was almost everywhere zero, making it difficult to apply the backpropagation algorithm. According to (cybenko89), for a neural network to satisfy the universal approximation property, the activation function must be bounded, non-constant, monotonically increasing, and continuous. In the 1980s and 1990s, smooth activations like the sigmoid activation and tanh activation, along with their variations, were proposed (Rumelhart86, Narayan97, LeCun02). These activations had the disadvantage of gradient vanishing when stacking layers deeply. To overcome this drawback, the Rectified Linear Unit (ReLU) was proposed by (Nair10), whose derivative is a non-zero constant. Variations of ReLU, such as PReLU, GELU, and ELU, have also been proposed in the 2010s (He15, Hendrycks16, Clervert15). Recently, there have been attempts to find activations tailored to specific tasks to impart inductive bias to the neural network structure according to the task (Ramachandran17, Sun24).
@@ 1.2 Rectified Power Unit
One of the attempts to provide good regularity and inductive bias to the architecture is the Rectified Power Unit (RePU), a generalization of ReLU (Li20). Early results using RePU include the application of ReQU, a type of RePU, in the DeepRitz method, which is a neural network approach for solving Partial Differential Equations (PDEs) (E18). RePU has also been proposed to construct differentiable neural networks, and error bounds related to this have been analyzed (Li20). According to research, using a RePU network allows for the optimal approximation of smooth functions in Sobolev space with optimal depth and width, and the optimal approximation error bound is also provided in Li20. Similarly, another study derived that a shallow RePU network can approximate multivariate functions of a certain function class with finite neural network weight norms and possibly unbounded width (Abdeljawad22). Furthermore, it was shown that the derivatives of a network composed of RePU can be expressed as a mixture of various types of RePU, and an improved upper bound on the complexity of these derivatives was derived. This demonstrates the capability of RePU by showing the risk bound for the deep score matching estimator (DSME) (Shen23). Despite these results highlighting the good regularity of RePU neural networks, experimental evidence shows that RePU networks suffer from exploding or vanishing values and fail to train properly when the layers are deep.
1.3 Effective Field Theory of Neural Networks
The language of quantum field theory (QFT) is useful for analyzing systems with a very high degree of freedom. One such example is artificial neural networks. It is known that the distribution of neural network ensembles follows a Gaussian process when the width is infinite, corresponding to the free field in the QFT framework (Neal96, Rasmussen04, Roberts22, Banta24, Halverson21). In practice, neural networks with finite width exhibit non-Gaussianity in their distributions as layers are stacked, allowing the distribution of neurons to be calculated perturbatively, as in weakly-interacting theory (Banta24, Halverson21). Interestingly, when the width of the neural network is finite, the propagation of information through the neural network can be understood as a renormalization group (RG) flow. By analyzing the recursive relations of the kernels in RG flow, hyperparameters can be tuned to a criticality state that ensures the stability of the network (Roberts22). Additionally, this analysis allows for the classification of activation functions into universality classes based on the behavior of the RG flow (Roberts22). Some types of activation functions, according to this classification, cannot satisfy criticality conditions from the outset.
@@ 1.4 Our Contribution
This paper provides a comprehensive analysis of the Rectified Power Unit (RePU) activation function and its limitations through the lens of effective field theory. Based on these insights, we propose the Modified Rectified Power Unit (MRePU) activation function, which addresses the drawbacks of RePU while preserving its advantages. Our contributions are summarized as follows:
  • Theoretical Analysis of RePU’s Limitations: We derive the susceptibility of RePU activations and demonstrate that they fail to satisfy criticality conditions, leading to instability in deep neural networks. This analysis identifies why RePU networks experience vanishing or exploding kernels during forward propagation, irrespective of hyperparameter initialization.
  • Proposal of MRePU: Based on our theoretical insights, we propose the Modified Rectified Power Unit (MRePU) activation function. MRePU retains the differentiability and universal approximation properties of RePU while overcoming its instability issues. We show that MRePU belongs to a distinct universality class, satisfying criticality conditions necessary for stable training.
  • Approximation Properties of MRePU Networks: We theoretically prove that MRePU networks possess universal approximation properties for differentiable functions, akin to RePU networks. Furthermore, MRePU networks exhibit efficient polynomial approximation.
  • Experimental Validation and Inductive Bias: Extensive experiments validate the theoretical predictions, demonstrating that MRePU introduces a specific inductive bias suitable for tasks requiring smooth approximations and accurate differentiation. While MRePU does not universally outperform ReLU or GELU across all tasks, it shows superior performance in approximating derivatives and achieving stable learning in tasks involving polynomial and differentiable function approximation. For example, MRePU networks provide meaningful improvements in training stability and accuracy for deep architectures, including those applied to physics-informed neural networks (PINNs).
  • Proposal of Criticality Condition for Hyperparameters and Empirical Validation with Real-World Tasks: Based on our theoretical findings, we derived the criticality condition required for the statistical distribution of the initial ensemble of MRePU networks to ensure stable training. We empirically verified, by constructing phase diagrams, that this condition serves as the determinant boundary between training success and failure. Furthermore, we confirmed the effective performance of MRePU networks on real-world tasks, such as MNIST and CIFAR-10, employing widely adopted deep architectures including ResNet.
@@ 2 Preliminary
In this section, we define the neural network and activation functions that will be the subject of our analysis, focusing particularly on the Rectified Power Unit (RePU). Additionally, we explore the effective field theory of neural networks, which will be the primary tool for our analysis.
@@ 2.1 Overview of Neural Networks and Activation Functions
Definition 1 (Fully Connected Network (FCN)). A Fully Connected Network $f(x;\theta):\mathbb{R}^{n_{in}}\rightarrow\mathbb{R}^{n_{out}}$ is defined by the following recursive equations:
@@
Definition 2 (Initialization distribution of biases and weights). In the definition of neural networks, there are adjustable weights and biases parameters. The method of setting these parameters at the beginning of training is called initialization, and typically, each parameter follows a specific probability distribution, referred to as the initialization distribution. Although it is possible to use mathematically complex initialization distributions, it is common practice to assign each weight and bias under the condition of being independent and identically distributed (i.i.d). The probability distributions used can include Gaussian distribution, truncated Gaussian distribution, uniform distribution, and others. For wide neural networks, it is known that when the mean of the weight distribution is 0 and the variance is constant, differences arising from different types of distributions are suppressed by $\frac{1}{width}$. Therefore, for the sake of convenience in our discussion, we will set the distributions of weights and biases to follow a mean-zero Gaussian distribution:
@@
[…]
Sigmoid $\sigma(z)=\frac{1}{1+e^{-z}}$ cybenko89
Tanh $\sigma(z)=\tanh(z)=\frac{e^{z}-e^{-z}}{e^{z}+e^{-z}}$ cybenko89, LeCun02
Sin $\sigma(z)=\sin(z)$ Parascandolo16, Sitzmann20
[…]
Softplus $\sigma(z)=\log(1+e^{z})$ Bishop06
SWISH $\sigma(z)=\frac{z}{1+e^{-z}}$ Ramachandran17
GELU $\sigma(z)=\Big[\frac{1}{2}+\frac{1}{2}\text{erf}\Big(\frac{z}{\sqrt{2}}\Big)\Big]z$ Hendrycks16
@@
Definition 3 (Rectified Power Unit (RePU). The RePU activation function is defined by the following equation:
@@
where $p$ is a positive integer that determines the power to which the input $z$ is raised. When $p=1$, it corresponds to ReLU, and thus RePU can be considered a generalized version of ReLU. The cases where $p=2$ and $p=3$ are referred to as Rectified Quadratic Unit (ReQU) and Rectified Cubic Unit (ReCU), respectively. For RePU with $p=k$, it is easy to see that it is differentiable k1k-1 times, and this property ensures that the neural network and its gradients are differentiable functions.
@@ 2.2 Effective Field Theory for Neural Networks
Definition 4 ($M$-point correlators). The moments or $M$-point correlators of a probability distribution $p(z)=p(z_{1},...,z_{N})$ are defined as follows:
$\mathbb{E}[z_{\mu_{1}}\cdots z_{\mu_{M}}]=\int z_{\mu_{1}}\cdots z_{\mu_{M}}p(z)d\nu(z).$
Definition 5 ($M$-point connected correlators). For $2\leq M$, the cumulant or $M$-point connected correlators of a probability distribution $p(z)$ are defined as follows:
Proposition 6 (Wick contraction). Suppose $(X_{1},\dots,X_{n})$ is a zero-mean multivariate normal random vector. Then, all odd-order correlators vanish, and the even-order correlators are given by
$\mathbb{E}[X_{j_{1}}^{i_{1}}\dots X_{j_{n}}^{i_{n}}]=\sum_{\text{(all possible parings)}}\prod_{\text{(parings)}}\mathbb{E}[X_{p_{1}}X_{p_{2}}]$
For $M=1$, the connected correlator coincides with the regular correlator. For parity-symmetric distributions, it also coincides for $M=2$. Using Wick contraction, it is known that for a Gaussian distribution, the connected correlators for $M>2$ are zero. Therefore, higher-point connected correlators can be used to determine how much the distribution deviates from a Gaussian distribution. Additionally, if the connected correlators for $M>2$ are small, the distribution is defined as nearly-Gaussian. In quantum field theory, the 2-point connected correlator can be seen as an expression of the translational symmetry of a free field. In a free field, there is no interaction due to potential, so all connected correlators higher than the 2-point vanish. Conversely, if there are interactions between particles due to a potential, higher-order connected correlators become non-zero. Similarly, in the neural network we analyze, if the higher-order correlators do not vanish and instead show significant values, this implies meaningful interactions between nodes in the feature space, suggesting that feature learning is taking place.
Induced distributions Following the notation in (1), let $p(z^{(l)}|\mathcal{D})$ denote the probability distribution of the preactivations in the $l$-th layer given a dataset $\mathcal{D}=\{x_{i;\alpha}\}_{i=1,...,n_{0};\alpha=1,...,N_{\mathcal{D}}}$. Here, $\alpha$ denotes the label index for the samples. For the first layer, the preactivation has $M>2$ connected correlators equal to 0, and has the following mean and covariance, indicating that each preactivation follows an independent Gaussian distribution.
$\begin{split}\mathbb{E}[z_{i;\alpha}^{(1)}]&=\mathbb{E}\bigg[b_{i}^{(1)}+\sum_{j=0}^{n_{0}}W_{ij}^{(1)}x_{j;\alpha}\bigg]=0,\\ \mathbb{E}[z_{i_{1};\alpha_{1}}^{(1)}z_{i_{2};\alpha_{2}}^{(1)}]&=\mathbb{E}\bigg[\bigg(b_{i_{1}}^{(1)}+\sum_{j=0}^{n_{0}}W_{i_{1}j_{1}}^{(1)}x_{j_{1};\alpha_{1}}\bigg)\bigg(b_{i_{2}}^{(1)}+\sum_{j=0}^{n_{0}}W_{i_{2}j_{2}}^{(1)}x_{j_{2};\alpha_{2}}\bigg)\bigg]\\ &=\delta_{i_{1}i_{2}}\bigg(C_{b}^{(1)}+C_{W}^{(1)}\frac{1}{n_{0}}\sum_{j=1}^{n_{0}}x_{j;\alpha_{1}}x_{j;\alpha_{2}}\bigg)=\delta_{i_{1}i_{2}}G_{\alpha_{1}\alpha_{2}}^{(1)}.\end{split}$
@@
$\begin{split}&S(z^{(l)}):=\frac{1}{2}\sum_{i=1}^{n_{l}}\sum_{\alpha_{1}\alpha_{2}\in\mathcal{D}}g^{\alpha_{1}\alpha_{2}}_{(l)}z_{i;\alpha_{1}}^{(l)}z_{i;\alpha_{2}}^{(l)}\\ &-\frac{1}{8}\sum_{i_{1},i_{2}=1}^{n_{l}}\sum_{\alpha_{1},...,\alpha_{4}\in\mathcal{D}}v^{(\alpha_{1}\alpha_{2})(\alpha_{3}\alpha_{4})}_{(l)}z_{i;\alpha_{1}}^{(l)}z_{i;\alpha_{2}}^{(l)}z_{i;\alpha_{3}}^{(l)}z_{i;\alpha_{4}}^{(l)}+O((z^{(l)})^{6}).\end{split}$
It is known that the following relationship holds for the coefficients. This can be derived by comparing the coefficients of the connected correlator given by the mean metric and four-point vertex with those of the connected correlator assumed by the ansatz provided by action $S$. $g^{\alpha_{1}\alpha_{2}}_{(l+1)}$ and $v^{(\alpha_{1}\alpha_{2})(\alpha_{3}\alpha_{4})}_{(l+1)}$ (Roberts22):
$\begin{split}g^{\alpha_{1}\alpha_{2}}_{(l)}&=G^{\alpha_{1}\alpha_{2}}_{(l)}+O\Big(\frac{1}{n_{l-1}}\Big),\\ v^{(\alpha_{1}\alpha_{2})(\alpha_{3}\alpha_{4})}_{(l)}&=\frac{1}{n_{l}}V^{(\alpha_{1}\alpha_{2})(\alpha_{3}\alpha_{4})}_{(l)}+O\Big(\frac{1}{n_{l-1}^{2}}\Big).\end{split}$
@@
Kernels and Representation Group flow Now, we will investigate how the statistics of the preactivation change as they pass through the layers. We will refer to the evolution of the distribution through the layers as the representation group flow, following the terminology of (Roberts22). While it is not feasible to explicitly calculate the statistics for a complex non-Gaussian distribution, it is possible to develop a perturbative theory to understand the recursive flow of the key statistics of the distribution. Before delving into this, let’s define the notation for Gaussian expectations. For an observable $\mathcal{O}(z)$ with a mean-zero Gaussian distribution with variance $\bf{K}=(K_{ij})$, the expectation is denoted as $\langle\mathcal{O}(z)\rangle_{\bf{K}}$. Now, if we define the kernel of the neural network through the following recursive relation:
$\begin{split}K^{(1)}_{\alpha_{1}\alpha_{2}}&:=G^{(1)}_{\alpha_{1}\alpha_{2}},\\ K^{(l+1)}_{\alpha_{1}\alpha_{2}}&=C^{(l+1)}_{b}+C^{(l+1)}_{W}\langle\sigma(z_{\alpha_{1}})\sigma(z_{\alpha_{2}})\rangle_{K^{(l)}}.\end{split}$ (4)
it is known that this kernel satisfies following properties, as discussed in (Roberts22).
$\begin{split}G^{(l)}_{\alpha_{1}\alpha_{2}}&=K^{(l)}_{\alpha_{1}\alpha_{2}}+O\Big(\frac{1}{n_{l-1}}\Big),\\ V^{(l+1)}_{(\alpha_{1}\alpha_{2})(\alpha_{3}\alpha_{4})}&=\Big(C_{W}^{(l+1)}\Big)^{2}\Big[\langle\sigma(z_{\alpha_{1}})\sigma(z_{\alpha_{2}})\sigma(z_{\alpha_{3}})\sigma(z_{\alpha_{4}})\rangle_{K^{(l)}}\\ &-\langle\sigma(z_{\alpha_{1}})\sigma(z_{\alpha_{2}})\rangle_{K^{(l)}}-\langle\sigma(z_{\alpha_{3}})\sigma(z_{\alpha_{4}})\rangle_{K^{(l)}}\\ &+\frac{n_{l}}{4n_{l-1}}\sum_{\beta_{1},...,\beta_{4}\in\mathcal{D}}V^{(\alpha_{1}\alpha_{2})(\alpha_{3}\alpha_{4})}_{(l)}\Big\langle\sigma(z_{\alpha_{1}})\sigma(z_{\alpha_{2}})\Big(z_{\beta_{1}}z_{\beta_{2}}-K_{\beta_{1}\beta_{2}}^{(l)}\Big)\Big\rangle_{K^{(l)}}\\ &\times\Big\langle\sigma(z_{\alpha_{3}})\sigma(z_{\alpha_{4}})\Big(z_{\beta_{3}}z_{\beta_{4}}-K_{\beta_{3}\beta_{4}}^{(l)}\Big)\Big\rangle_{K^{(l)}}\Big]+O\Big(\frac{1}{n_{l}}\Big).\end{split}$ (5)
@@
Then, it is known that the following recursion can be derived (Roberts22):
@@
Summary of Framework In this section, we investigate the evolution of preactivation statistics across layers, referred to as the Representation Group flow. By analyzing the recursive kernel relation (Eq. (4)), we derive a fixed point for the statistical flow. When the initial ensemble of the network operates near this fixed point, the propagation of small perturbations through the layers is governed by the parallel ($\mathcal{X}_{\parallel}$) and perpendicular ($\mathcal{X}_{\perp}$) susceptibilities. These quantities can be interpreted as the eigenvalues of the linearized flow along the respective directions around the fixed point. Consequently, for both training dynamics and inference, it is crucial that these eigenvalues are close to unity ($\mathcal{X}\approx 1$). This criticality condition ensures that perturbations are neither amplified nor suppressed but are preserved across layers, thereby maintaining a stable signal propagation.